OpenStudy (anonymous):

Use a graphing calculator to solve the equation -3 cos t = 1 in the interval from 0 to 2p. Round to the nearest hundredth.

OpenStudy (anonymous):

I typed it in on my calculator but did not do a thingg!

OpenStudy (swissgirl):

\(\cos(t)=\large{1 \over- 3} \) \(\cos^{-1}(\large{1 \over- 3} )=t\)

OpenStudy (anonymous):

How did you get that?

OpenStudy (swissgirl):

First I divided -3 from both sides

OpenStudy (swissgirl):

\(-3\cos(t)=1\) \(\large {-3\cos(t) \over -3}={1 \over -3}\) \(\cos(t)=\large{1 \over -3}\) \(\cos^{-1}(\cos(t))=Cos^{-1}(\large{1 \over -3})\) \(t=Cos^{-1}(\large{1 \over -3})\)

OpenStudy (anonymous):

How would you put this in a decimal form.?

OpenStudy (swissgirl):

lol you gotta put this into your calc first and then your answer will be a decimal

OpenStudy (anonymous):

Oh wow that was dumb of me lol. I was thinking this had to be put in a decimal form.! before solving it

OpenStudy (anonymous):

I got 109.47

OpenStudy (swissgirl):

Are you supossed to be working in radians or degrees?

OpenStudy (anonymous):

Ah, I had my calc set in degrees, I'm pretty sure I'm suppose to be using radians.! Gah I keep forgetting to change the mode

OpenStudy (anonymous):

Ok now I got 1.91.!

OpenStudy (swissgirl):

YUP THAT IS WHAT I GOT :)

OpenStudy (anonymous):

Thanks for the help.! (-: